OpenStudy (superhelp101):

Using complete sentences, explain how you would factor completely x9 - 64.

OpenStudy (superhelp101):

@jdoe0001

OpenStudy (broskishelleh):

x³ - y³ = (x - y)(x² + xy + y²) (x³)³ - (4)³ = (x³ - 4)(x^6 + 4x³ + 16)

OpenStudy (broskishelleh):

Amirite?

OpenStudy (broskishelleh):

A medal y'all!

OpenStudy (superhelp101):

what's amirite?

OpenStudy (jdoe0001):

\(\bf x^9 - 64\implies (x^3)^3-4^3\quad and\implies {\color{brown}{ a^3-b^3 = (a-b)(a^2+ab+b^2)}}\)

OpenStudy (broskishelleh):

Has anyone here done 9.06 in Algebra 1

OpenStudy (superhelp101):

not taking it

OpenStudy (broskishelleh):

And it means am I right?

OpenStudy (superhelp101):

oh lol. Thx guys! :)

OpenStudy (jdoe0001):

yw